Fred can repair the roof in 4 hours. If Sandy helps, it takes them 3 hours. Without help, how long would it take for Sandy to repair the roof?
Answer by richwmiller(17219)

rate for A alone 4
rate for A & B together 3
1/a+1/b=1/x
1/4+1/b=1/3
1/b=1/3-1/4
1/b=4/12-3/12
1/b=1/12
1/b=1/12
b=12
rate for B alone 12
check
1/4+1/12=1/3
12/48+4/48=1/3
16/48=1/3
ok
codewa